You recently purchased a stock that is expected to earn 19 percent in a booming economy, 8 percent in a normal economy, and lose 28 percent in a recessionary economy. There is a 20 percent probability of a boom and a 70 percent chance of a normal economy. What is standard deviation on this stock?
Answer #1

A B C D E F 1 ER Prob 19% 2 3 4 5 Booming Normal Recession Deviation from mean 0.0030752 0.0001372 0.0119716 20% 70% 10% 8% -

А B с D E 1 2 ER 3 Booming Normal Recession 0.19 0.08 Prob 0.2 0.7 = 1-03-04 Deviation from mean =(C3-$D$8)^2*D3 =(C4-$D$8)^2